# Uncertainty-Aware Measurement of Scenario Suite Representativeness for Autonomous Systems

### v1

@@ IV-B Problem Setting: Scenario Suite
The scenario suite was constructed to serve as an empirical test distribution against which the TOD prior could be compared. Its purpose is to emulate the process of autonomous-vehicle scenario generation, where rare but safety-critical conditions are intentionally oversampled to ensure sufficient coverage for evaluation.
@@
where each conditional term reflects both the dependency structure and the applied weighting multipliers. After normalization, the resulting probabilities formed a valid joint distribution:
@@ IV-C Problem Setting: TOD
with posterior mean per category which is used for illustration purpose.
@@
Evaluating this expression at $n^{(0)}=\underline{n}$ and $n^{(0)}=\overline{n}$ yields the lower and upper posterior expectations, defining credible bounds on the TOD distribution. In the numerical examples we have applied $n^{(0)}=10$.
@@ IV-D Numerical results
Figure 2: Comparison between the empirical scenario suite distribution and the posterior mean of the TOD distribution.
@@
The near overlap between the orange (suite) and blue (posterior TOD) points indicates a strong alignment for most categories. Small deviations, typically for adverse or nighttime conditions, reflect deliberate over-sampling in the scenario suite design. This visualization helps identify which operational categories contribute most to representativeness discrepancies before summarizing them with global metrics such as the TVD or JSD.
@@
As the prior strength $n^{(0)}$ increases, the posterior distribution becomes more influenced by the prior and less by the observed scenario frequencies, leading to a gradual increase in $D_{\mathrm{TV}}$ VI. This monotonicity caused by iLUCK model [imprecision_luck_walter_2009] reflects how epistemic confidence in the prior shifts the posterior away from the empirical scenario suite (small width of the TVD intervals, e.g., $[0.0031,0.0045]$, for small width of $n^{(0)}$, e.g., $n^{(0)}\in[15,20]$).
@@ V Discussion
V-A Prior distribution
@@
In addition, published reference statistics—such as national traffic surveys, meteorological archives, or the multi-source environmental datasets used in Hiller et al. [hiller_creating_2022]—can provide baseline frequency estimates for relevant factors like weather type, road category, or traffic density. When no credible information exists, a uniform prior $\boldsymbol{y}^{(0)}=(1/K,\ldots,1/K)$ expresses a state of complete ignorance by assigning equal probability to all possible categories.
V-B Dependence and Independence among TOD Categories
In this study, the TOD is defined as a combination of discrete operational factors such as weather, road type, illumination, and traffic density etc.
To derive the joint distribution, it is necessary to account for the dependencies that may exist among these factors, as such factors are rarely independent in real-world operational environments. Certain weather conditions may be more frequent on rural roads than in urban areas; night-time driving is typically associated with lower traffic density; and fog or heavy rain may be confined to specific geographical regions or seasons. These dependencies imply that the true TOD distribution $\boldsymbol{\theta}_{\mathrm{TOD}}$ cannot, in general, be expressed as a simple product of marginal probabilities:
@@ V-B Dependence and Independence among TOD Categories
Neglecting such correlations can lead to systematic under- or over-representation of certain operational categories within the scenario suite, potentially biasing the assessment of representativeness.
In the presented example (Sec. IV), rather than assuming independence among these factors, we explicitly assumed dependencies to better reflect realistic operational conditions. These dependencies need to be determined based on prior expert knowledge and supported by available historical data, which indicate systematic relationships between certain environmental and contextual factors. If the dependencies are not well understood and independence between variables must be assumed, a sensitivity analysis can help to avoid misleading conclusions by quantifying the potential impact of this assumption on the representativeness metrics.
@@ VI Summary and Conclusions
This paper presented a probabilistic framework for assessing the representativeness of scenario datasets with respect to the Target Operational Domain (TOD). Building upon the definition of representativeness in ISO 8800 [ISO8800_2024], the proposed approach models the TOD as an uncertain categorical distribution over joint operational conditions and applies a robust Bayesian formulation with imprecise Dirichlet priors to capture epistemic uncertainty about both prior beliefs and data sufficiency. The resulting posterior credal set enables transparent quantification of uncertainty in representativeness metrics—specifically the Total Variation Distance (TVD) and Jensen–Shannon Divergence (JSD)—between the observed scenario suite distribution and the inferred TOD distribution. This probabilistic treatment extends the existing approaches by embedding them within a formal Bayesian uncertainty quantification framework.
The numerical illustration demonstrated how the method distinguishes between systematic dataset biases (e.g., intentional over-sampling of safety-critical scenarios) and epistemic uncertainty due to limited or imprecise knowledge of operational frequencies. The interval-valued discrepancy measures provide interpretable and auditable evidence for safety assurance arguments, aligning with current developments in trustworthy AI and scenario-based testing.
Future work will extend the current framework along several directions. First, we will address the dependence between the TOD and scenario suite distributions, acknowledging that both often stem from the same limited real-world dataset. Treating them as fully independent can conceal shared biases and limit the ability to reveal representational gaps. Next, we aim to relax the simplifying assumptions made in this study by incorporating continuous operational factors, thereby moving beyond purely discretized representations toward a hybrid framework that combines discrete and continuous variables. Beyond representativeness, we will develop quantitative metrics for other data-related safety properties—such as completeness, consistency, and relevance—and explore their interrelations within a unified framework. Ultimately, the goal is to embed these quantitative assessments within a complete Assurance 2.0 case, linking probabilistic evidence to structured, auditable safety arguments for trustworthy AI systems.
Appendix A Outline of the Proof of Theorem 1
@@
Hence,
$\boldsymbol{\theta}_{\text{TOD}}\mid\mathbf{k}\;\sim\;\mathrm{Dirichlet}\!\big(n^{(0)}\boldsymbol{y}^{(0)}+\mathbf{k}\big),$
and each component marginally follows:
$\theta_{k}\!\mid\!\mathbf{k}\sim\mathrm{Beta}(a_{k},a_{0}-a_{k})$
with $a_{k}=n^{(0)}y^{(0)}_{k}+k_{k}$ and $a_{0}=n^{(0)}+n$.
@@
$\Delta_{K-1}$. Under the iLUCK [imprecision_luck_walter_2009] monotonicity result, the lower and upper bounds of this joint posterior family (the posterior envelope) are attained at the extreme points of $\mathcal{Y}_{\text{TOD}}\times[\underline{n},\overline{n}]$. Monotonicity holds because, under the iLUCK reparameterization, the Dirichlet shape parameters

### v2

@@ IV-B Problem Setting: Scenario Suite
The scenario suite serves as an empirical test distribution for comparison with the inferred TOD distribution, emulating autonomous-vehicle scenario generation where rare but safety-critical conditions are intentionally oversampled to ensure sufficient evaluation coverage.
@@
where each conditional term reflects both the dependency structure and applied weighting multipliers. After normalization, the probabilities form a valid joint distribution:
@@
As shown in Fig. 1, the numerical example illustrates three distributions: the empirical scenario-suite distribution obtained directly from the 1600 generated scenarios; the unknown true TOD distribution representing real operational exposure; and the imprecise TOD distribution, given by the interval-valued posterior inferred from limited operational samples and prior assumptions. Representativeness is assessed by comparing the scenario-suite distribution with this uncertain TOD distribution.
@@ IV-C Problem Setting: TOD
with posterior mean per category (for illustration purpose):
@@
Evaluating this expression at $n^{(0)}=\underline{n}$ and $n^{(0)}=\overline{n}$ yields the lower and upper posterior expectations, defining credible bounds on the TOD distribution.
@@ IV-D Numerical results
Figure 2: Comparison between the empirical scenario suite distribution and the posterior mean of the TOD distribution (assuming $n^{(0)}=10$).
@@
The near overlap between the orange (suite) and blue (posterior TOD) points indicates a strong alignment for most categories. Small deviations, typically for adverse or nighttime conditions, reflect deliberate over-sampling in the scenario suite design. This visualization highlights which operational categories contribute most to representativeness discrepancies before summarization with global metrics (e.g., TVD, JSD).
A high discrepancy indicates that the scenario suite does not reflect real operational conditions. Per-category differences can identify under- and over-represented situations, guide scenario generation or reduction, and iteratively refine the suite until the discrepancy falls below an acceptable level.
@@
As the prior strength $n^{(0)}$ increases, the posterior distribution becomes more influenced by the prior and less by the observed scenario frequencies, leading to a gradual increase in $D_{\mathrm{TV}}$ VI. This monotonicity caused by iLUCK model [12] reflects how epistemic confidence in the prior shifts the posterior away from the empirical scenario suite (small width of the TVD intervals, e.g., $[0.0031,0.0045]$, for small width of $n^{(0)}$, e.g., $n^{(0)}\in[15,20]$).
While the numerical example considers a scenario suite broadly aligned with the inferred TOD, the proposed metrics also capture strong mismatch: when scenarios are generated independently of the TOD structure or uniformly at random, TVD and JSD increase toward their upper bounds, indicating poor representativeness and enabling identification of highly non-representative suites.
@@ V Discussion
V-A Limitations of single-valued representativeness metrics
As discussed in Section II, existing approaches assume the reference distribution is known with sufficient confidence. In data-sparse settings, this is problematic: single-valued representativeness metrics conceal sensitivity to prior assumptions, whereas interval-valued results reveal the range of plausible conclusions. These are not meant to replace single-valued metrics, but to complement them when data are limited or uncertain by exposing epistemic uncertainty that single-point results would otherwise hide.
V-B Prior distribution
@@
In addition, published reference statistics—such as national traffic surveys, meteorological archives, or the multi-source environmental datasets used in Hiller et al. [7]—can provide baseline frequency estimates for relevant factors like weather type, road category, or traffic density. When no credible information exists, a uniform prior $\boldsymbol{y}^{(0)}=(1/K,\ldots,1/K)$ expresses a state of complete ignorance by assigning equal probability to all possible categories.
V-C Dependence and Independence among TOD Categories
In this study, the TOD is defined as a combination of discrete operational factors such as weather, road type, illumination, and traffic density etc. To derive the joint distribution, it is necessary to account for the dependencies that may exist among these factors, as such factors are rarely independent in real-world operational environments. Certain weather conditions may be more frequent on rural roads than in urban areas; night-time driving is typically associated with lower traffic density; and fog or heavy rain may be confined to specific geographical regions or seasons. These dependencies imply that the true TOD distribution $\boldsymbol{\theta}_{\mathrm{TOD}}$ cannot, in general, be expressed as a simple product of marginal probabilities:
@@ V-C Dependence and Independence among TOD Categories
Neglecting such correlations can lead to systematic under- or over-representation of certain operational categories in the scenario suite, biasing the representativeness assessment.
In Sec. IV, dependencies were explicitly modeled to reflect realistic operational conditions. These dependencies need to be determined based on prior expert knowledge and supported by available historical data, which indicate systematic relationships between certain environmental and contextual factors. If the dependencies are not well understood and independence between variables must be assumed, a sensitivity analysis can help to avoid misleading conclusions by quantifying the potential impact of this assumption on the representativeness metrics.
@@ VI Summary and Conclusions
This paper presented a probabilistic framework for assessing the representativeness of scenario datasets with respect to the TOD. Unlike existing approaches based on fixed distributions and single-point discrepancy measures, the proposed approach models the TOD as an uncertain categorical distribution over joint operational conditions and applies a robust Bayesian formulation with imprecise Dirichlet priors to capture epistemic uncertainty about both prior beliefs and data sufficiency. The resulting posterior credal set enables transparent quantification of uncertainty in representativeness metrics (e.g., Total Variation Distance) between the observed scenario suite distribution and the inferred TOD distribution.
Unlike the numerical example, real-world ODDs involve many variables; the framework therefore relies on structured or factorized representations capturing only relevant dependencies.
Future work will extend the current framework along several directions. First, we will relax the i.i.d. assumption to account for potential correlations between scenarios, which may arise from temporal, spatial, or operational dependencies in real-world driving data. Next, we will address the dependence between the TOD and scenario suite distributions, acknowledging that both often stem from the same limited real-world dataset. Treating them as fully independent can conceal shared biases and limit the ability to reveal representational gaps. We also aim to relax the simplifying assumptions made in this study by incorporating continuous operational factors, thereby moving beyond purely discretized representations toward a hybrid framework that combines discrete and continuous variables. Beyond representativeness, we will develop quantitative metrics for other data-related safety properties—such as completeness, consistency, and relevance—and explore their interrelations within a unified framework. Ultimately, these assessments should be embedded within an Assurance 2.0 case (e.g., [3]), linking probabilistic evidence to structured, auditable safety arguments for trustworthy AI. While the numerical example demonstrates the feasibility and interpretability of the approach, broader empirical comparisons with existing representativeness metrics on larger and more realistic datasets are left for future work.
Acknowledgment
This work is supported by the European Union’s EU Framework Program for Research and Innovation Europe Horizon (grant agreement No 101202457) and Innovate UK through the DriveSafeAI Project (10062594). XZ’s contribution is supported by the UK EPSRC New Investigator Award [EP/Z536568/1]. SK’s contribution is supported by the UKRI Future Leaders Fellowship Grant [MR/S035176/1]. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Proof Sketch of Theorem 1
@@
Hence, $\boldsymbol{\theta}_{\text{TOD}}\mid\mathbf{k}\;\sim\;\mathrm{Dirichlet}\!\big(n^{(0)}\boldsymbol{y}^{(0)}+\mathbf{k}\big),$ and each component marginally follows: $\theta_{k}\!\mid\!\mathbf{k}\sim\mathrm{Beta}(a_{k},a_{0}-a_{k})$ with $a_{k}=n^{(0)}y^{(0)}_{k}+k_{k}$ and $a_{0}=n^{(0)}+n$.
@@
$\Delta_{K-1}$. Under the iLUCK [12] monotonicity result, the lower and upper bounds of this joint posterior family (the posterior envelope) are attained at the extreme points of $\mathcal{Y}_{\text{TOD}}\times[\underline{n},\overline{n}]$. Monotonicity holds because, under the iLUCK reparameterization, the Dirichlet shape parameters